# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ A.3.3. Satisfying Conditions (I) to (III)
which is a contradiction when $\epsilon>0$, which implies that condition (III) holds.
Appendix B Overview of [20]
For $\epsilon\geq 0$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $\epsilon$-DP IM w.r.t. a neighboring relation $\sim$. Suppose that $\mathcal{M}$ has discrete answer distributions. Vadhan and Wang [20] shows that for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ are identical. In this section, we overview the construction of $\mathcal{P}$ and show it satisfies the following property:
Claim B.1.
For any $t\in\mathbb{N}$ and any query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, if the distributions of the first $t$ answers of $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to the queries $(q_{1},\dots,q_{t})$ are identical, then for every $b\in\{0,1\}$, when $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,0}(b)$ receives the queries $q_{1},\dots,q_{t}$, $\mathcal{P}$ generates answers itself, without interacting with $\mathrm{RR}_{\epsilon,0}(b)$.
Since $\delta=0$, $\mathrm{RR}_{\epsilon,0}(b)$ returns $(b,\bot)$ with probability $\frac{e^{\epsilon}}{1+e^{\epsilon}}$ and $(1-b,\bot)$ otherwise. For each $b$ and query sequence $(q_{1},\dots,q_{t})\in Q_{\mathcal{M}}^{t}$, let $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$ denote the distribution of the first $t$ answers of $\mathcal{M}(s_{b})$ in response to $q_{1},\dots,q_{t}$. If $\epsilon=0$, then define $\nu_{q_{1},\dots,q_{t}}^{b}$ to be identical to $\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}$. Otherwise, set
$\nu_{q_{1},\dots,q_{t}}^{b}=\frac{e^{\epsilon}\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{b})}-\mu_{q_{1},\dots,q_{t}}^{\mathcal{M}(s_{1-b})}}{e^{\epsilon}-1}$
For $t\in\mathbb{N}$, $b\in\{0,1\}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q_{\mathcal{M}}\times A_{\mathcal{M}})^{t-1}\times Q$, if $\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})>0$, the function $T_{q_{1},\dots,a_{t-1},q_{t}}:A_{\mathcal{M}}\to[0,1]$ is defined as
$T_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=\frac{\nu_{q_{1},\dots,q_{t}}^{b}(q_{1},\dots,a_{t})}{\nu_{q_{1},\dots,q_{t-1}}^{b}(q_{1},\dots,a_{t-1})}$
Vadhan and Wang [20] define $\mathcal{P}$ as follows: The initial state of $\mathcal{P}$ is $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence. Let $h$ denote the sequence. Let $(r,\bot)$ denote the outcome of $\mathrm{RR}_{\epsilon,0}(b)$. After interacting with $\mathrm{RR}_{\epsilon,0}(b)$, $\mathcal{P}$ updates null to $r$. Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$. To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the distribution $T_{h,q}^{r}$, where $h$ is the history of all previous queries and answers.
[…]

### v3

@@ 8.3.4. Satisfying Conditions (I) to (IV)
which is a contradiction when $\epsilon>0$, which implies that condition (III) holds. ∎
To show $\phi_{t}$ satisfies conditions (IV), we need the following lemma:
Lemma 8.32.
For any query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then
$\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$.
Proof.
Fix $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$. We will inductively show that for every $a_{t}\in A$,
$\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}.$
Recall $\xi_{t}(q_{1},\dots,a_{t})$ is defined as the lexicographically maximum feasible solution $(x_{0}^{*},x_{1}^{*})$. By Constraint ($\xi$.4), for each $b\in\{0,1\}$, we have
$x_{b}^{*}\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}.$
Therefore, to prove the lemma, it suffices to show
$x_{0}=x_{1}=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
is a feasible solution:
  • ($\xi$.1): By Lemma 8.17 and non-negativity of $\xi_{t}$, this constraint is satisfied.
  • ($\xi$.2): To show Constraint ($\xi$.2) is satisfied, we need to show for each $b\in\{0,1\}$,
    $\sum_{a\in\Gamma_{<a_{t}}\cup\{a_{t}\}}$ $\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\}$
    $+\sum_{a\in\Gamma_{>a_{t}}}\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
    Since
    $\mathrm{L}_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a)\leq\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\},$
    it suffices to prove
    $\sum_{a\in A}$ $\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\}\leq\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})$
    By condition (IV) and Remark 8.21,
    $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\min\{\mu_{t-1}^{b}(q_{1},\dots,a_{t-1}),\mathrm{L}_{t-1}^{0}(q_{1},\dots,a_{t-1})+\mathrm{L}_{t-1}^{1}(q_{1},\dots,a_{t-1})\}$
    Therefore, by Lemma 8.10 and Lemma 8.20,
    $\phi_{t-1}^{b}(q_{1},\dots,a_{t-1})\geq\min\{\mu_{t-1}^{b}(q_{1},\dots,a_{t-1}),\mathrm{L}_{t-1}^{0}(q_{1},\dots,a_{t-1})+\mathrm{L}_{t-1}^{1}(q_{1},\dots,a_{t-1})\}$
    $\geq\min\left\{\sum_{a\in A}\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\sum_{a\in A}\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\sum_{a\in A}\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\right\}$
    $\geq\sum_{a\in A}\min\{\mu_{t}^{b}(q_{1},\dots,a_{t-1},q_{t},a),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t-1},q_{t},a)+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t-1},q_{t},a)\},$
    completing the proof.
  • ($\xi$.3): Recall that $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$. To prove this constraint is satisfied, we need to show
    $(1-e^{-\epsilon})\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}\leq(1-e^{-\epsilon})\mu_{t}^{0}(q_{1},\dots,a_{t}),$
    which is trivially true.
  • ($\xi$.4): This constraint is trivially satisfied as equality holds.
Claim 8.33.
$\phi_{t}$ satisfies conditions (IV).
Proof.
For any query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if
($\star$) $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$
for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then by Lemma 8.32,
($\star\star$) $\xi_{t}^{0}(q_{1},\dots,a_{t})=\xi_{t}^{1}(q_{1},\dots,a_{t})=\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
for all $(a_{1},\dots,a_{t})\in A^{t}$. Moreover, by Lemma 8.10 and $\star$8.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$. Thus, by Remark 8.21 and condition (IV),
($\star\star\star$) $\phi_{t-1}^{0}(q_{1},\dots,a_{t-1})=\phi_{t-1}^{1}(q_{1},\dots,a_{t-1}).$
Therefore, by equations ($\star$8.3.4), ($\star\star$8.3.4) and ($\star\star\star$8.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$. Hence, there exists a maximal feasible solution $(z^{*},z^{*})$ satisfying all the constraints. By definition of $\phi_{t}(q_{1},\dots,a_{t})$, we have $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$. Also, by Constraint ($\phi$.1), $z^{*}$ is at least $\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$.
9. $f^{\delta}_{\mathrm{RR}}$-concurrent composition
In this section, we prove:
Lemma (Restatement of Lemma 5.9).
For every $0\leq\delta\leq 1$, the $f^{\delta}_{\mathrm{RR}}$-concurrent composition of continual mechanisms is $(0,\delta)$-differentially private.
In Section 4, we assumed that all composed CMs have discrete answer distributions. Accordingly, we also prove Lemma 5.9 under this assumption. During this section, we assume the domain of all variables is finite or countable. For every random variable $Y$, we denote its support by $\mathrm{supp}(Y)$.
We begin by proving some intermediate results needed for the final proof. Vadhan et al. [24] show that to guarantee an IM is differentially private, it suffices to consider only deterministic adversaries. Specifically, the following holds:
Lemma 9.1 ([24]).
An IM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a neighboring relation $\sim$ if and and only if for every deterministic adversary $\mathcal{A}$ and every pair of neighboring initial states $s\sim s^{\prime}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{M}(s))$ and $\mathrm{View}(\mathcal{A},\mathcal{M}(s^{\prime}))$ are $(\epsilon,\delta)$-indistinguishable.
By Definition 3.7, this result extends naturally to CMs:
Corollary 9.2.
A CM $\mathcal{M}$ is $(\epsilon,\delta)$-DP w.r.t. a verification function $f$ if and and only if for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!\mathcal{M})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!\mathcal{M})$ are $(\epsilon,\delta)$-indistinguishable.
The definition of $(0,\delta)$-DP CMs can be equivalently expressed using the notion of total variation:
Definition 9.3 (Total Variation).
Let $Y^{0}$ and $Y^{1}$ be two variables with the same domain $\mathcal{Y}$. The total variation distance between $Y^{0}$ and $Y^{1}$, denoted $\mathrm{TV}(Y^{0},Y^{1})$, is defined as
$\mathrm{TV}(Y^{0},Y^{1})=\sup_{S\subseteq\mathcal{Y}}\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S].$
Lemma 9.4.
Let $Y^{0}$ and $Y^{1}$ be two random variables over the same domain $\mathcal{Y}$. For every $0\leq\delta\leq 1$, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if $\mathrm{TV}(Y^{0},Y^{1})\leq\delta$.
Proof.
By definition, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if for every $S\subseteq\mathcal{Y}$, $\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S]\leq\delta$. Therefore, $Y^{0}$ and $Y^{1}$ are $(0,\delta)$-indistinguishable if and only if $\mathrm{TV}(Y^{0},Y^{1})=\sup_{S\subseteq\mathcal{Y}}\Pr[Y^{0}\in S]-\Pr[Y^{1}\in S]\leq\delta$. ∎
Lemma 9.5.
[…]
$\forall k>0,\quad\mathrm{TV}\left((W_{1}^{0},\dots,W_{k}^{0}),(W_{1}^{1},\dots,W_{k}^{1})\right)\leq\delta.$
Then,
$\mathrm{TV}(V^{0},V^{1})\leq\delta.$
Proof.
We have
$\mathrm{TV}(V^{0},V^{1})$ $=\frac{1}{2}\sum_{\sigma\in\Sigma^{*}}\big|\Pr[V^{0}=\sigma]-\Pr[V^{1}=\sigma]\big|$
$=\frac{1}{2}\sum_{\sigma\in\Sigma^{*}}\big|\Pr[W^{0}=\sigma\cdot\mathdollar^{\omega}]-\Pr[W^{1}=\sigma\cdot\mathdollar^{\omega}]\big|$
$=\frac{1}{2}\sum_{\sigma^{\prime}\in\Sigma^{*}\cdot\mathdollar^{\omega}}\big|\Pr[W^{0}=\sigma^{\prime}]-\Pr[W^{1}=\sigma^{\prime}]\big|=\mathrm{TV}(W^{0},W^{1})$
Thus, we need to show $\mathrm{TV}(W^{0},W^{1})\leq\delta$.
For each $i\in\mathbb{N}$ and $b\in\{0,1\}$, let $W_{1:i}^{b}=(W_{1}^{b},\dots,W_{i}^{b})$ denote the prefix of length $i$, and let $W_{i:}^{b}=(W_{i}^{b},W_{i+1}^{b},\dots)$ denote the infinite suffix of $W^{b}$. Define:
$T=\mathrm{TV}(W^{0},W^{1}),\quad T_{i}=\mathrm{TV}(W_{1:i}^{0},W_{1:i}^{1}),\quad\text{and}\quad\alpha(i)=\Pr[W_{i}^{0}\neq\mathdollar]+\Pr[W_{i}^{1}\neq\mathdollar$].
We will show that
  1. (1)
    $T_{i}$ is non-decreasing in $i$.
  2. (2)
    For all $i\in\mathbb{N}$, we have $T-T_{i}\leq\alpha(i+1)$.
Since each $T_{i}\leq 1$ and the sequence is non-decreasing, the limit $T^{*}=\lim_{i\to\infty}T_{i}$ exists. Moreover, since $W_{i}^{0}=\mathdollar$ if and only if $|V^{b}|<i$ and $V^{b}$ is finite with probability $1$, the probability $\Pr[W_{i}^{b}\neq\mathdollar$] converges to zero as $i\to\infty$. Therefore,
$T-T^{*}=\lim_{i\to\infty}T-T_{i}\leq\lim_{i\to\infty}\alpha(i)=0.$
Since $T_{i}\leq\delta$ for all $i$, we conclude that $T^{*}\leq\delta$. Thus, $T=\mathrm{TV}(W^{0},W^{1})\leq\delta$.
It remains to prove (1) and (2). Since $W_{1:i}^{0}$ and $W_{1:i}^{1}$ in $T_{i}=\mathrm{TV}(W_{1:i}^{0},W_{1:i}^{1})$ equal $W_{1:i+1}^{0}$ and $W_{1:i+1}^{1}$ in $T_{i}=\mathrm{TV}(W_{1:i+1}^{0},W_{1:i+1}^{1})$ with their last entries removed, the data-processing inequality implies $T_{i}\leq T_{i+1}$. For (2), define $X_{i}^{b}=W_{1:i}^{b}\mathdollar^{\omega}$ for each $b\in\{0,1\}$. By definition, $T_{i}=\mathrm{TV}(X_{i}^{0},X_{i}^{1})$. We have
$T-T_{i}$ $=\mathrm{TV}(W^{0},W^{1})-\mathrm{TV}(X_{i}^{0},X_{i}^{1})$
$=\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[W^{1}\in S]\right)-\sup_{S}\left(\Pr[X^{0}\in S]-\Pr[X^{1}\in S]\right)$
$\leq\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[W^{1}\in S]-\Pr[X^{0}\in S]+\Pr[X^{1}\in S]\right)$
$\leq\sup_{S}\left(\Pr[W^{0}\in S]-\Pr[X^{0}\in S]\right)+\sup_{S}\left(\Pr[X^{1}\in S]-\Pr[W^{1}\in S]\right)$
$=\Pr[W^{0}\neq X_{i}^{0}]+\Pr[W^{1}\neq X_{i}^{1}]$
[…]
$T-T_{i}\leq\Pr[W_{i+1}^{0}\neq\mathdollar]+\Pr[W_{i+1}^{1}\neq\mathdollar]=\alpha(i+1),$
finishing the proof. ∎
Lemma 9.6 (The Maximal Coupling Lemma).
Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$. There exists a joint distribution for a pair of random variables $(Z^{0},Z^{1})$ on $\mathcal{Y}\times\mathcal{Y}$, called a coupling, such that:
  1. (i)
    For every $b\in\{0,1\}$, the marginal distribution of $Z^{b}$ is identical to the distribution of $Y^{b}$.
  2. (ii)
    Let $S^{0}=\mathrm{supp}(Y^{0})$ and $S^{1}=\mathrm{supp}(Y^{1})$. The probability of the event $Z^{0}=Z^{1}$ is maximized and satisfies
    $\Pr[Z^{0}=Z^{1}]=\sum_{y\in\mathcal{Y}}\min\left\{\Pr[Z^{0}=y],\Pr[Z^{1}=y]\right\}=\sum_{y\in S^{0}\cap S^{1}}\min\left\{\Pr[Z^{0}=y],\Pr[Z^{1}=y]\right\}.$
Lemma 9.7 (The Coupling Lemma).
Let $Y^{0}$ and $Y^{1}$ be random variables with domain $\mathcal{Y}$. For every (not necessarily maximal) coupling $(Z^{0},Z^{1})$ of $(Y^{0},Y^{1})$, where the marginals of $Z^{0}$ and $Z^{1}$ are identically distributed as of $Y^{0}$ and $Y^{1}$, we have
$\mathrm{TV}(Y^{0},Y^{1})\leq\Pr[Z^{0}\neq Z^{1}].$
Lemma 9.8.
Let $k\in\mathbb{N}$, and let $Y^{0}=(Y_{1}^{0},\dots,Y_{k}^{0})$ and $Y^{1}=(Y_{1}^{1},\dots,Y_{k}^{1})$ be random variables. For $i\in[k-1]$ and $b\in\{0,1\}$, set $S_{i}^{b}=\mathrm{supp}(W_{1:i}^{b})$ and $S_{i}=S_{i}^{0}\cap S_{i}^{1}$, with $S_{0}=S_{0}^{b}=\emptyset$. For $i\in[k]$, let $\Delta_{i}:S_{i-1}\to[0,1]$ be a function, and fix $\gamma\in[0,1]$. Suppose the following conditions hold:
  • For every $i\in[k]$ and every $y_{1:i-1}\in S_{i-1}$,
    $\mathrm{TV}\left((Y_{i}^{0}|Y_{1:i-1}^{0}=y_{1:i-1}),(Y_{i}^{1}|Y_{1:i-1}^{1}=y_{1:i-1})\right)\leq\Delta_{i}(w_{1:i-1}).$
  • For every $y_{1:k-1}\in S_{k-1}$, $1-\prod_{j=1}^{i}\left(1-\Delta_{j}(w_{1:j})\right)\leq\gamma$.
Then,
$\mathrm{TV}(Y^{0},Y^{1})\leq\gamma.$
Proof.
We construct a coupling $\left(Z^{0}=(Z_{1}^{0},\dots,Z_{k}^{0}),Z^{1}=(Z_{1}^{1},\dots,Z_{k}^{1})\right)$ for $(Y^{0},Y^{1})$ and show that $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$. By Lemma 9.7, this implies $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. Throughout the proof, let $\mathcal{Y}$ denote the union of all possible outcomes of $Y_{i}^{b}$ over $i\in[k]$ and $b\in\{0,1\}$.
Define $(Z_{1}^{0},Z_{1}^{1})$ to be the maximal coupling of $Y_{1}^{0}$ and $Y_{1}^{1}$ given by Lemma 9.6. For $i=2$ to $k$, given samples $(Z_{1:i-1}^{0},Z_{1:i-1}^{1})$, we define $(Z_{i}^{0},Z_{i}^{1})$ as follows:
  • If the histories are identical, i.e. $Z_{1:i-1}^{0}=Z_{1:i-1}^{1}=y_{1:i-1}$, then $(Z_{i}^{0},Z_{i}^{1})$ is drawn from the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}$}, given by Lemma 9.6.
  • Otherwise, if $Z_{1:i-1}^{0}=y_{1:i-1}^{0}$ and $Z_{1:i-1}^{1}=y_{1:i-1}^{1}$ differ, then $Z_{i}^{0}$ and $Z_{i}^{1}$ are drawn independently from their respective conditional distributions $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}^{0}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}^{1}$}.
By the chain rule, the probability $\Pr[Z^{0}=Z^{1}]=\sum_{y_{1:k}\in S_{k}}\Pr[Z^{0}=Z^{1}\!=\!y_{1:k}]$ equals
$\sum_{y_{1:k}\in S_{k}}\prod_{i=1}^{k}\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]=$
$\sum_{y_{1:k-1}\in S_{k-1}}\prod_{i=1}^{k-1}\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]\!\!\!\!\sum_{\begin{subarray}{c}y_{k}\in\mathcal{Y}\\ \text{s.t. }y_{1:k}\in S_{k}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k}^{0}\!=\!Z_{k}^{1}\!=\!y_{k}\,\middle|\,Z_{1:k-1}^{0}\!=\!Z_{1:k-1}^{1}\!=\!y_{1:k-1}\right].$
Conditioning on $Z_{1:k-1}^{0}=Z_{1:k-1}^{1}=y_{1:k-1}$, the pair $(Z_{k}^{0},Z_{k}^{1})$ is the maximal coupling of $Y_{i}^{0}\mid Y_{1:i-1}^{0}=y_{1:i-1}^{0}$} and $Y_{i}^{1}\mid Y_{1:i-1}^{1}=y_{1:i-1}^{1}$}. The innermost sum is taken over the common support of these conditional random variables. Thus, by Lemma 9.6, this summation equals
$1-\mathrm{TV}\left((Y_{k}^{0}\mid Y_{1:k-1}^{0}=y_{1:k-1}),(Y_{k}^{1}\mid Y_{1:k-1}^{1}=y_{1:k-1})\right).$
Hence, by the first assumption,
(6) $\sum_{\begin{subarray}{c}y_{k}\in\mathcal{Y}\\ \text{s.t. }y_{1:k}\in S_{k}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k}^{0}\!=\!Z_{k}^{1}\!=\!y_{k}\,\middle|\,Z_{1:k-1}^{0}\!=\!Z_{1:k-1}^{1}\!=\!y_{1:k-1}\right]\geq 1-\Delta_{k}(y_{1:k-1})$
By the second assumption, we know
$1-\Delta_{k}(y_{1:k-1})\geq\frac{1-\gamma}{\prod_{i=1}^{k-1}(1-\Delta_{i}(y_{1:i-1}))}.$
Therefore,
$\Pr[Z^{0}=Z^{1}]$ $\geq(1-\gamma)\sum_{y_{1:k-1}\in S_{k-1}}\prod_{i=1}^{k-1}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
$=(1-\gamma)\sum_{y_{1:k-2}\in S_{k-2}}\prod_{i=1}^{k-2}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
$\qquad\qquad\qquad\qquad\quad\sum_{\begin{subarray}{c}y_{k-1}\in\mathcal{Y}\\ \text{s.t. }y_{1:k-1}\in S_{k-1}\end{subarray}}\!\!\!\!\frac{\Pr\!\left[Z_{k-1}^{0}\!=\!Z_{k-1}^{1}\!=\!y_{k-1}\,\middle|\,Z_{1:k-2}^{0}\!=\!Z_{1:k-2}^{1}\!=\!y_{1:k-1}\right]}{1-\Delta_{k-1}(y_{1:k-2})}.$
The term $\Delta_{k-1}(y_{1:k-2})$ is independent of $y_{k-1}$. Moreover, by an argument identical to the one for Inequality 6, we have
$\sum_{\begin{subarray}{c}y_{k-1}\in\mathcal{Y}\\ \text{s.t. }y_{1:k-1}\in S_{k-1}\end{subarray}}\!\!\!\!\Pr\!\left[Z_{k-1}^{0}\!=\!Z_{k-1}^{1}\!=\!y_{k-1}\,\middle|\,Z_{1:k-2}^{0}\!=\!Z_{1:k-2}^{1}\!=\!y_{1:k-1}\right]\geq 1-\Delta_{k-1}(y_{1:k-2}).$
Thus,
$\Pr[Z^{0}=Z^{1}]\geq(1-\gamma)\sum_{y_{1:k-2}\in S_{k-2}}\prod_{i=1}^{k-2}\frac{\Pr\!\left[Z_{i}^{0}\!=\!Z_{i}^{1}\!=\!y_{i}\,\middle|\,Z_{1:i-1}^{0}\!=\!Z_{1:i-1}^{1}\!=\!y_{1:i-1}\right]}{1-\Delta_{i}(y_{1:i-1})}$
A backward iteration with the same proof implies that $\Pr[Z^{0}=Z^{1}]\geq 1-\gamma$. Consequently, $\Pr[Z^{0}\neq Z^{1}]\leq\gamma$, and by Lemma 9.7, $\mathrm{TV}(Y^{0},Y^{1})\leq\gamma$. ∎
Proof of Lemma 5.9.
By Corollary 9.2 and the definition of concurrent composition using the continual mechanism ExtConComp, it suffices if we show that for every deterministic adversary $\mathcal{A}$, the random variables $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(0)\!\circ^{*}\!{\textrm{ExtConComp}})$ and $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(1)\!\circ^{*}\!{\textrm{ExtConComp}})$ are $(0,\delta)$-indistinguishable. Fix a deterministic adversary $\mathcal{A}$.
Since $\mathcal{A}$ is deterministic, its sequence of random coins is empty (or predetermined), and its $i$-th query message to the verifier $\mathcal{V}[f^{\delta}_{\mathrm{RR}}]$ can be determined by the first i1i-1 responses from the verifier. Therefore, there exists a post-processing function $g$ such that for every $b\in\{0,1\}$, $g(V^{b})$ is equivalently distributed as $\mathrm{View}(\mathcal{A},\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}})$, where $V^{b}$ is the (finite) sequence of responses $V^{b}_{i}$ sent by $\mathcal{V}[f^{\delta}_{\mathrm{RR}}]\!\circ^{*}\!\mathcal{I}(b)\!\circ^{*}\!{\textrm{ExtConComp}}$ to $\mathcal{A}$. By the (non-interactive) post-processing lemma, it suffices to prove that the random variables $V^{0}$ and $V^{1}$ are $(0,\delta)$-indistinguishable.
[…]
As $\mathcal{A}$ is deterministic, for every $i\in\mathbb{N}$ and $w_{1:i-1}=(w_{1},\dots,w_{i-1})\in S_{i-1}^{b}$, the random variable $W_{i}^{b}$ conditioned on $W_{1:i-1}=w_{1:i-1}$ is distributed as follows:
  • If $w_{i-1}=\mathdollar$ (i.e., the communication has halted before) or the design of $\mathcal{A}$ is such that it halts the communication after receiving the i1i-1-st response $w_{i-1}$ given the history $w_{1:i-1}$, then $W_{1:i-1}=\mathdollar$ with probability $1$.
  • Otherwise, $W_{1:i-1}$ is equivalently distributed to $\mathrm{RR}_{0,\delta_{i}^{\prime}}(b)$, where $\delta_{i}^{\prime}=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ is determined by the history $w_{1:i-1}$ and the design of $\mathcal{A}$.
Define $\Delta_{i}(w_{1:i-1})=0$ in the former case and $\Delta_{i}(w_{1:i-1})=\delta_{i}^{\prime}(w_{1:i-1},\mathcal{A})$ in the latter. By Lemma 9.4 and Lemma 2.4, for every $i\in\mathbb{N}$ and every $w_{1:i-1}\in S_{i-1}^{b}$,
$\mathrm{TV}\left((W_{i}^{0}|W_{1:i-1}^{0}=w_{1:i-1}),(W_{i}^{1}|W_{1:i-1}^{1}=w_{1:i-1})\right)\leq\Delta_{i}(w_{1:i-1}).$
[…]